Shift Summary for Evening shift, 11/19/99

Patricia McBride (mcbride@fnpsph)
Sat, 20 Nov 1999 00:11:56 CST

Shift Summary for Evening shift, 11/19/99

On Shift: Mike Arenton, Patty McBride

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 7.9
Hours of E799 data (special runs not included): 6.9
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: The hunt for the hyperon trigger problem, one swap
of an Etot controller board, a rad trip after an access, and one
dart_fresh_start.

========================================================================
Runs ended this shift:

Run 15013 17.3 M BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift:

Sydney and Patty made an access to swap the Etot controller board for the
bottom east quadrant

========================================================================
Overview of shift:
E799 data taking with one access to replace Etot board.

A problem with DC4x hole R DCOR rate created a hyperon trigger problem (L1/sem
alarms for trigger 10, 11). The DC4X H R rate looks fine at the end of the
cable from downstairs. We tried kfrends, reloading STT, L1, L2 and a cold
start of the trigger. the problem went away
a few minutes later, but returned just before midnight. Jason and Leo take
charge and set out to
change the trigger configuration.

========================================================================
Description of Shift:

run spill time comment
- ------------------------------------------------------------------------
15013 15:30 L1_B10/NM2SEM, L1_B11/NM2SEM errors on every spill.
We start to investigate and find a problem with the
x vertex distribution from the hyp799 kumac. There are
almost no events with x less than 0 ;
We page Leo, Rick Kessler, Eric Ramberg

16:00 Eric suggests that it is a DCOR problem so we try to
find Ashkan and Theo. They are not available, so we
page Hogan.

16:20 Beam off for 15 minutes

312 HCC inefficiency error; reload ETOT to clear a problem
column in
the bottom east quadrant

Hogan, Leo, Arun, BoBT, Ashkan (by phone), Mike and
Patty
investigate the trigger and DCOR scaler rates and
find a problem with the DC4x_H_R scaler rates. This
is also
evident in the trg timing plots.

Leo and Ashkan work on a file that would change the
hole trigger
so that it will not use DC4 (old configuration).

316 Cold start of the trigger

330 the trigger problem goes away... so we stick with the
trigger
configuration with DC4x for now.
Another HCC inefficiency warning, reload ETot maps

334 Reload Etot maps again
358 Reload Etot maps again, reload HCC maps too

20:10 Sydney changes the compression ratio in the dis
database.
Tapes should now reach 95% before the run ends.

384 HCC inefficiency again, same place; reload Etot maps
400 reload Etot maps
410 reload Etot maps
417 reload Etot maps
425 21:00 call for an access to replace Etot controller board
for the
bottom east quadrant

21:25 Out of access; do a dart_fresh_start so that
compression ratio
change will take effect

15017 20:50 triggers once again; things look okay, so we call and
ask for beam

22:10 CAll MCR, they tell us about a Rad trip; they are
investigating

24 22:18 Beam arrives

~97 23:55 Alarm for L1_B10/NM2SEM, L1_B11/NM2SEM sounds again.

99 24:00 Cold start the trigger. Problem persists. Leo
phones...
hand off to new shift crew

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Label 30 UPH tapes if bored on shift.


-- 
Patricia McBride          mcbride@fnal.gov          PHONE: (630)840-8071
Fermi National Accelerator Laboratory               FAX:   (630)840-2783